### v1

# Lost in Translation, Found in Context: Sign Language Translation with Contextual Cues

Abstract
Our objective is to translate continuous sign language into spoken language text. Inspired by the way human interpreters rely on context for accurate translation, we incorporate additional contextual cues together with the signing video, into a new translation framework. Specifically, besides visual sign recognition features that encode the input video, we integrate complementary textual information from (i) captions describing the background show, (ii) translation of previous sentences, as well as (iii) pseudo-glosses transcribing the signing. These are automatically extracted and inputted along with the visual features to a pre-trained large language model (LLM), which we fine-tune to generate spoken language translations in text form. Through extensive ablation studies, we show the positive contribution of each input cue to the translation performance. We train and evaluate our approach on BOBSL – the largest British Sign Language dataset currently available. We show that our contextual approach significantly enhances the quality of the translations compared to previously reported results on BOBSL, and also to state-of-the-art methods that we implement as baselines. Furthermore, we demonstrate the generality of our approach by applying it also to How2Sign, an American Sign Language dataset, and achieve competitive results.
1 Introduction
Sign languages are the natural means of communication for deaf communities [67]. They are visual-spatial languages and lack standardised written forms [26, 3]. Sign languages exist independently of spoken languages, with their own unique lexicons, ordering, and grammatical structures. Furthermore, sign languages are expressed through both manual and non-manual components, e.g. mouthings, which can be conveyed simultaneously [74, 19].
The focus of this work is sign language translation (SLT), the process of transforming sign language into spoken language, in the open-vocabulary setting. Achieving SLT could significantly improve accessibility and inclusion for the deaf and hard-of-hearing communities by reducing communication barriers.
A key challenge in SLT is the fact that sign languages rely heavily on contextual discourse and spatial awareness due to their visual-spatial nature [53, 68, 49, 43]. Indeed, a study by [68] with fluent deaf signers found that a third of their sentence-level signing clips could only be fully translated when provided with additional discourse-level context. We next describe three examples of contextual dependencies in sign language: (a) Signers use spatial indexing to identify referents introduced earlier in the discourse (e.g. pointing to identify a particular person, object or any previously defined concept placed in signing space) [44, 25, 4]; (b) Sign languages typically follow a topic-comment structure [67]. For instance, the topic can refer to setting the temporal framework (e.g. if ‘yesterday’ is signed, all verbs in the translation should be in the past tense until a new time is mentioned). Once a topic is setup, it is usually not mentioned at every sentence but only re-established when changed, making a sentence-level translation often ill-defined; (c) Finally, sign language exhibits homonyms [67], where two signs with similar hand movements can have different meanings (e.g. ‘battery’ and ‘uncle’ in British Sign Language (BSL)).
Another major obstacle to automatic SLT is the scarcity of large-scale training data. Sign languages are low resource languages, with limited availability of signing videos online, and manually translating signing content is extremely time-consuming, requiring expert annotators.
In this paper, we turn to the underexplored and large-scale BOBSL dataset [2], which consists of over 1,400 hours of BSL interpreted TV broadcasts with accompanying English subtitles, as our source of training data. In comparison to datasets with sentence-level translations [11, 29, 85], interpreters in BOBSL translate spoken language subtitles to signing using context, both from the video playing in the background as well as the previous discourse that has been signed (see Fig. 1). We leverage this setting to explore the benefit of context for SLT performance. Specifically, we use background descriptions from a captioning model and predicted translations of previous sentences, along with sign-level pseudo-glosses111We abuse the linguistic gloss term and refer to sign-level translations in free-form English as glosses. and strong signing visual features as inputs to a pre-trained LLM, which we fine-tune to generate spoken language translations in text form. As signers may use pointing to identify an object or person on the screen, the background descriptions can help with identifying the corresponding referents (e.g. there are 4,179 pointing occurrences in the BOBSL-CSLR test set annotations [60], which spans only 6 hours). Furthermore, ambiguities of homonyms and tenses/co-references due to a topic-comment structure may be resolved thanks to the context provided from previous sentences, as well as the background.
However, this task remains challenging due to the weak and noisy nature of the TV subtitle supervision. The supervision is weak because the subtitles are temporally aligned with the speech, and not perfectly with the signing. Although we employ existing automatic signing-subtitle alignment methods [7], subtitles may appear a few seconds before or after the corresponding signing sequence. Additionally, the supervision is noisy because words in the subtitle are not necessarily signed, and vice versa. Indeed, sign language interpretation corresponds to a translation – as opposed to a transcription – of the speech content, and can often lead to simplification in vocabulary [6]. Furthermore, during TV broadcasts, where time pressure is a factor, interpreters may occasionally omit content to keep up with the spoken audio.
In this work, we make the following contributions: (i) we propose a new LLM-based model that integrates visual signing and text features with contextual information, including video background descriptions and previous sentence translations; (ii) we conduct a thorough analysis to examine the impact of each input cue on the translation quality, and introduce an LLM-based scoring mechanism that provides a more nuanced translation assessment than traditional metrics such as BLEU [56]; (iii) we evaluate previous state-of-the-art models [75, 84] on the BOBSL dataset to establish a performance benchmark and find our proposed model surpasses them significantly; (iv) we show our contextual method generalises to How2Sign, an American Sign Language (ASL) dataset, demonstrating its effectiveness.
2 Related Work
We discuss relevant works on isolated and continuous sign language recognition, as well as sign language translation.
Isolated Sign Language Recognition (ISLR), which involves classifying a short single-sign video clip into a sign category, has been extensively researched over many years. In the past decade, ISLR has made significant strides, largely due to the emergence of deep spatiotemporal neural networks and the availability of larger-scale datasets [1, 50, 9, 8, 34, 40, 41, 37]. In particular, the I3D model [13] has demonstrated its effectiveness in providing robust features for recognition [37, 40, 1]. More recently, the Video Swin Transformer [48] has shown strong ISLR results, and been employed as a backbone for various sign language tasks involving continuous video streams [57, 60]. In this work, we leverage the sign video encoder from [60] both to obtain visual features and pseudo-glosses.
Continuous Sign Language Recognition (CSLR) typically involves recognising gloss sequences – spoken language words corresponding to semantic labels for individual signs – from a continuous signing video clip. It is crucial to distinguish CSLR from translation due to the significant grammatical differences between signed and spoken languages [67]. Benchmarks such as PHOENIX14 [39] and CSL-Daily [85], which have manually glossed training data, are widely used for CSLR. Due to lack of temporal alignment between signs and video frames, many studies [20, 18, 73, 35, 10, 87, 30, 58, 51, 83] utilise the Connectionist Temporal Classification (CTC) loss [27]. More recently, [60] demonstrates the advantages of using video-to-text retrieval on the challenging BOBSL dataset [2].
Automatic Sign Language Translation (SLT) involves converting sign language videos into spoken language. Given that glosses act as a mid-level representation that bridge the visual signing and spoken language modalities, CSLR has been employed as an intermediate step for sign language translation (known as Sign2Gloss2Text), or for pre/joint training to enhance visual representations [11, 12, 16, 85, 15, 80, 78, 81, 86, 77]. In our work, we only use sign-level pseudo-glosses as an additional input – rather than as auxiliary supervision – for translation.
Given the restricted scalability of manual glosses, recent research explores the gloss-free SLT setting [76, 53], utilising larger datasets and enhanced pre-training techniques. For instance, [2] and [69] train SLT systems from scratch using large-scale signing data and robust visual features from a pre-trained ISLR I3D backbone. [64] employs a similar approach, but further pre-trains the visual encoder for sign spotting. [46] trains an SLT system jointly with the visual backbone, leveraging conceptual anchor words. GFSLT [84] pre-trains their visual encoder and text decoder through a CLIP-style contrastive loss [59] and a masked self-supervised loss. [79] introduces a frame-wise contrastive loss during visual pre-training to enhance feature discrimination. VAP [36] pre-trains for more fine-grained visual and textual token alignment. [82] collects large-scale noisy multilingual YouTube SL data and jointly pre-trains for various tasks such as SLT, subtitle-signing alignment, and text-to-text translation. Similarly to these works, we turn to large-scale data and extract strong visual features from an ISLR model.
Recent works also incorporate large-scale language foundation models. For example, [66] inputs pseudo-glosses directly into ChatGPT. [70, 62, 61, 36, 82] all fine-tune a pre-trained T5 model for SLT. [70] feeds in 3D landmark embeddings, while [62] and [61] use visual features from a BEVT-pre-trained [72] ISLR model and a MAE-pretained video encoder, respectively. Sign2GPT [75] leverages large-scale pre-trained vision (DINOv2 [55]) and language (XGLM [47]) models, incorporating adapters (LoRA [32]) for transfer to sign language. Similarly to these works, we make use of a pre-trained LLM (Llama3 [24]) and fine-tune it for SLT.
The most closely related to our work is [65], that also makes use of context. Specifically, their approach encodes the ground truth previous subtitle, and spottings (automatic localised sign-level annotations obtained by querying words from the ground truth subtitle), as well as the signing video before passing them to a transformer decoder to generate translations. Our work differs on multiple fronts. Firstly, our method is fully automatic, employing the predicted translation of the previous sentence and pseudo-glosses from an ISLR model without access to any ground truth. Second, we additionally incorporate context from the background video. Lastly, we leverage a pre-trained LLM as opposed to training a language decoder from scratch.
3 Sign Language Translation with Context
We begin this section by outlining our framework for SLT using multifarious cues (Sec. 3.1). Next, we describe the representation of each of the inputs including visual features, pseudo-glosses, the background description, and the previous sentence (Sec. 3.2). We then present the training strategy that leverages a pre-trained LLM on this set of inputs (Sec. 3.3).
3.1 Framework overview
We introduce a new framework to perform SLT in an open-vocabulary setting, basing our model on a pre-trained LLM. As illustrated in Fig. 2, the proposed framework takes various cues as input: (i) visual features representing the signing video, (ii) pseudo-glosses as a (noisy) automatic transcription of signs, (iii) background description as a contextual cue from the TV show displayed behind the signer, and (iv) predicted translation of the previous sentence as another contextual cue, obtained auto-regressively at inference. All modalities except the visual features are in text form, which allows us to leverage the powerful language generation capabilities of an LLM, trained on diverse text corpora. To map the visual features into the space of the LLM input tokens, we train a simple MLP-based mapping network. Besides the above inputs, we provide the LLM with an instruction describing the SLT task: “You are an AI assistant designed to interpret a video of a sign language signing sequence and translate it into English.” Finally, we add single-sentence prompts in between the inputs by describing each input type, while serving as a separator (see Sec. A.2 for the exact prompts). All the inputs are appended into a single sequence and fed to the LLM.
3.2 Model inputs
In the following, we describe how we obtain each of the inputs.
Visual features. Following [60, 57], we employ a Video-Swin model [48] pre-trained for ISLR (classification of isolated signs) to obtain our visual features. Specifically, we utilise the recent and strong model provided by [60], which is trained on the BOBSL videos [2] for a vocabulary of 8,697 signs using the automatic sign spottings from [52]. This ISLR model processes short video clips of 16 frames (i.e. less than 1 second in 25 fps videos) to produce a single 768-dimensional embedding vector. Specifically, this vector comes from spatio-temporal averaging of the penultimate layer output of Video-Swin. To capture fine-grained temporal details, we feed 16 consecutive frames into the sign video encoder and apply a sliding window with a stride of to obtain features , where represents the number of visual features per sentence. For example, when =2, we have on average 56 features (i.e. 4.5 seconds) in BOBSL sentences. Note that for experiments conducted on the How2Sign dataset [22], we further fine-tune the Video-Swin model using the spotting annotations provided by [23] for a vocabulary of 1,887 signs (see Sec. A.5 for details).
Pseudo-glosses. To represent sign sequences in text form, we apply the ISLR model mentioned above in a sliding window manner to record the classification predictions, and obtain pseudo-glosses , representing a sequence of words (or phrases) from the vocabulary of the classifier (e.g. 8k signs). Note that these sign category predictions are noisy, often including more labels than the number of signs occurring in the sentence video, and many false positives, which we wish to suppress via our LLM tuning (see Fig. 2). Our pseudo-glosses are similar to [60], except we only apply repetition grouping, but we do not filter out low-confidence annotations with a threshold – this allows the LLM to learn which ones are relevant. In contrast to [60] that uses pseudo-glosses for supervision, we simply employ them as additional inputs to our SLT model. Typically, there are around 22 glosses per BOBSL sentence (which is less than the number of visual features – 56 features on average).
Background description. To incorporate context from the background footage, we crop out the signer from the full frame and apply an image captioning model (BLIP2 [42]) to extract textual descriptions of the scene behind the signer. Due to the repetitive nature of captions across consecutive frames and to reduce computational complexity, we extract captions at frames sampled at 1 fps, leaving us with 5 captions per signing sentence on average. Since similar scenes may persist even over several seconds, resulting in nearly identical captions, we collect all captions per sentence and keep only the list of unique words (e.g. 14 words per signing sentence). We further remove stopwords (as defined by [5]) to primarily feed keywords that may provide context, and consequently help disambiguate similar signs or identify pointings to the background screen. This process is illustrated in Sec. A.3.
Previous sentence. We incorporate the previous sentence text as an additional contextual cue. This refers to the sentence that the signer signed leading up to the current one. During training, we use either (i) the ground truth previous sentence or (ii) predictions precomputed from a model trained without the previous sentence cue (i.e. only visual features, pseudo-glosses, and background descriptions). At test time the model uses its own previous predictions as the previous sentence in an auto-regressive manner.
3.3 Tuning the LLM with multifarious cues
Given the inputs described above, we design and train an LLM-based model presented in the following.
Mapping network for visual features. All our inputs are already in text form except the visual features, which need a projection to map them into the text space, so that they can be fed into the pre-trained LLM. To this end, we train a simple mapping network, a 2-layer MLP with GELU [31] activation in between, projecting the visual features (of dimensionality 768) to the dimensionality of the LLM input embeddings (i.e. 4,096). We note that we add 1D temporal convolution layer to the mapping network when experimenting with the How2Sign dataset [22] (see Sec. A.5 for further details).
Training. The trainable parameters of our framework are in the mapping network and in the LLM. We randomly initialise the weights of the mapping network. For the LLM, we employ the open-source Llama3 model [24], specifically opting for the Llama3-8B variant to balance performance and efficiency. We tune the pre-trained LLM weights to adapt to our SLT task and to our input structure. Specifically, we adopt LoRA [32] fine-tuning (similarly to [75]) both to maintain computational efficiency, and not to degrade the powerful language decoding capability of the original model. We employ the standard cross-entropy loss across the original LLM vocabulary of 128k text tokens, using masked self-attention to predict the next tokens. At inference, the model auto-regressively decodes a sentence until an end-of-sentence token is reached.
Augmentations for textual modalities. During training, we apply several augmentations to our textual inputs to enhance model robustness. First, we perform word dropping on three textual cues (pseudo-glosses, the previous ground-truth sentence, and the background description) by randomly omitting between 0% and 50% of the words in each cue. Second, we randomly omit entire cues during training, each cue with 50% probability, allowing the model to flexibly handle inference when some modalities are missing, but also to make the model pay attention to each cue. Third, as previously mentioned, for previous sentences we make use of both ground-truth and predicted translations during training. Again, we randomly sample with a 50% probability between the two options. This strategy not only serves as an augmentation, but also reduces the domain gap between training and test time (where we do not have access to ground-truth previous sentences). In practice, we precompute the predicted previous sentences from a variant of our model trained with all cues except the previous sentence.
Implementation details. We train on 4 H100 GPUs with a batch size of 2 per GPU, utilising the Adam optimizer [38]. Training is performed in bfloat16 precision, with FlashAttention-2 [21] adapted to optimise memory usage. The LLM decoder (Llama3-8B model [24]) has a dimensionality 4,096 for its text embeddings. The LLM is fine-tuned using LoRA with a configuration of rank 4, alpha 16, and dropout 0.05. We fine-tune only the query and value projectors in all attention layers of the LLM. Since the text embedding layer has already been pre-trained on a large corpus, we freeze it during training. The training spans 10 epochs for BOBSL [2] and 15 epochs for How2Sign [22] datasets, including a warmup phase for the first 5 epochs with gradient clipping set to 1.0. The learning rate is set to 0.0001. We use the HuggingFace library for the pre-trained Llama3 models [24].
4 Experiments
In this section, we first present the datasets and a suite of evaluation protocols used in our experiments (Sec. 4.1), as well as baseline descriptions (Sec. 4.2). Next, we ablate various components of our framework (Sec. 4.3). We then show our improved translation performance compared to the state of the art on two challenging open-vocabulary benchmarks (Sec. 4.4). Finally, we illustrate qualitative results and discuss limitations (Sec. 4.5)
4.1 Data and evaluation protocols
BOBSL [2] comprises 1,500 hours of BSL-interpreted TV broadcast footage across a wide range of genres, accompanied by English subtitle sentences for the audio content. During training, to achieve better signing-sentence alignments, we use automatically signing-aligned sentences from [7] as described in [2]. We filter sentences to those lasting 1-20 seconds as in [60], resulting in 689k video-sentence training pairs, corresponding to a vocabulary of 86K words. For evaluation, we utilise the existing validation and test splits, Sent-Val and Sent-Test from [2], where English sentences have been manually aligned temporally to the continuous signing video. Sent-Val and Sent-Test consist of 1,973 and 20,870 aligned sentences, respectively, covering vocabularies of 3,528 and 13,641 English words. We report ablation studies on the validation set and present our final model results on the test set.
How2Sign [22] comprises 80 hours of ASL instructional videos from 10 different topics, with temporally aligned sentence language translations. There are 31,128 training, 1,741 validation and 2,322 test sentences, covering vocabularies of 15.7k, 3.2k, and 3.7k English words, respectively. We use the validation set to tune hyperparameters, and report our final model results on the test set.
Evaluation metrics. To evaluate translation performance effectively, we use five standard evaluation metrics: (i) BLEU-4 (B4) [56], which corresponds to the geometric mean of the precision scores of 4-grams, multiplied by a brevity penalty; (ii) BLEURT (B-RT) [63], which is a trained metric (using a regression model trained on ratings data) that can better capture non-trivial semantic similarities between sentences; (iii) ROUGE-L (R-L) [45], which measures the longest common subsequence between the prediction and ground truth sequence; (iv) CIDEr [71], a captioning metric that captures consensus of the prediction compared to ground truth by calculating the weighted cosine similarity of TD-IDF scores for various n-grams; and finally, (v) the Intersection over Union (IoU) of prediction and ground truth token sets (using the Penn Treebank tokenizer [5]). Similar to [60, 52], when computing IoU, we lemmatise all the words, and do not penalise translated words if they are synoyms to those in ground truth.
LLM Evaluation. Besides these standard metrics, we introduce an LLM-based evaluation metric adapted from the CLAIR framework [14] to assess sign language translations. We use the publicly available API of GPT-4o-mini from OpenAI [54], prompting the model to generate a score from 0 to 5 for each pair of translation and ground truth, where 5 indicates the best match and 0 is the worst, along with a detailed reasoning for the score. To calibrate the LLM, we provide 12 manually annotated in-context examples. We instruct the LLM to focus on key nouns and verbs and give less importance to pronouns. Further details are provided in Sec. A.1.
4.2 Baselines
Here, we review baselines that we compare against, including a common translation baseline (Albanie [2], Sincan [65]), as well as two state-of-the-art SLT methods (GFSLT [84], Sign2GPT [75]) according to PHOENIX14T [39, 11] and CSL-Daily [85] benchmarks.
Albanie [2] and Sincan [65]. These approaches build on the original SLT transformer work of Camgoz et al. [12], consisting of a standard transformer encoder-decoder architecture trained from scratch on pre-extracted video frame features. We compare against the first SLT baseline on BOBSL established by Albanie et al. [2] using this framework. Specifically, they train the I3D model for 2,281 sign vocabulary on BOBSL spottings, and supervise the translation with automatically-aligned English sentences filtered to a vocabulary of 9k common words. We also compare with the baseline of Sincan et al. [65], which similarly trains an encoder-decoder model and uses the same I3D features.
GFSLT [84]. This recent work introduces a pre-training phase featuring two components: (i) a CLIP-style contrastive loss, which directs the visual encoder to learn language-aligned visual representations, and (ii) a masked self-supervised loss, which promotes the ability of the text decoder to grasp sentence semantics. In the subsequent stage, the pre-trained visual encoder and text decoder are jointly fine-tuned within an encoder-decoder translation framework, enabling the direct conversion of visual representations into spoken sentences. We first reproduce this method using the public codebase on PHOENIX14T (see Sec. B.5), before adapting to the BOBSL dataset.
Sign2GPT [75]. This state-of-the-art work proposes an encoder-decoder translation framework that leverages large-scale pre-trained vision (DINOv2 [55]) and language (XGLM [47]) models, incorporating adapters (LoRA [32]) for transfer to sign language. Additionally, a prototype-driven pre-training strategy is introduced, which guides the visual encoder to learn sign representations from spoken language sentences by filtering specific parts of speech. We again use the public codebase for reproducing and applying on the BOBSL dataset. We report results both without and with pre-training (denoted as w/PGP) as in [75].
Oracle: Sincan [65] (Vid+Prev). This oracle baseline is a multi-modal variant of Sincan [65], where the three previous ground truth sentences are also fed as additional context to the decoder at both training and inference times.
Oracle: Sincan [65] (Vid+Prev +Spot). This baseline is another oracle variant, where Spottings are also fed into the decoder at both training and inference times. Spottings are automatic sign annotations [52], obtained using ground truth knowledge of the nearby English subtitles, i.e. given words from the subtitle, by temporally localising them in video. We note that our pseudo-glosses are different to Spottings as they are predicted directly from the video, without access to the corresponding English sentence translation.
4.3 Ablation study
In this section, we analyse our different design choices. We present our results on the BOBSL validation set, Sent-Val, using BLEURT (B-RT), IoU and LLM evaluation metrics.
Combining different cues. In Tab. 1, we measure the contribution of each input cue. With only visual features, the model achieves baseline scores of 41.0 for B-RT, 16.6 for IoU, and 1.29 for LLM score. Adding pseudo-glosses (i.e. textual cues derived from the current sign video) improves all metrics, highlighting the benefit of text directly related to the signs. Incorporating previous sentences as a contextual cue further boosts performance, and finally, adding background descriptions achieves the best results across all metrics. Overall, the final model, with all cues combined, yields a significant improvement of +2.5 in B-RT, +2.2 in IoU, and +0.27 in LLM score compared to the baseline, confirming that there is additional relevant information found in context, beyond the signing video, to help with the translation. In Sec. A.1, we provide further statistics analysing the distribution of the LLM evaluation scores, which align well with other metrics in most cases, while being interpretable.
| Vid | PG | Prev | BG | B-RT | IoU | LLM |
| ✓ | 41.0 | 16.6 | 1.29 | |||
| ✓ | ✓ | 41.8 | 17.5 | 1.40 | ||
| ✓ | ✓ | ✓ | 42.5 | 18.1 | 1.45 | |
| ✓ | ✓ | ✓ | ✓ | 43.5 | 18.8 | 1.56 |
| LLM | Drop | Drop | Prev/ | B-RT | IoU | LLM |
| fine-tuning | words | cue | Prev | |||
| ✓ | 41.2 | 17.0 | 1.40 | |||
| ✓ | ✓ | 41.4 | 17.4 | 1.41 | ||
| ✓ | ✓ | ✓ | 42.7 | 18.1 | 1.53 | |
| ✓ | ✓ | ✓ | ✓ | 43.5 | 18.8 | 1.56 |
| ✓ | ✓ | ✓ | 40.6 | 16.7 | 1.27 |
Effect of augmentations. We perform a series of ablations in Tab. 2 regarding the input augmentations at training time: (i) Drop words: randomly removing up to 50% of the words in textual cues, (ii) Drop cue: randomly removing an entire cue with a probability of 50%, and (iii) Prev/Prev: sampling either predicted or ground truth (GT) previous sentence with equal probability. When Drop words augmentation is applied, we observe a slight performance increase. Adding Drop cue augmentation provides an additional improvement (42.7 vs 41.4). We hypothesise that these augmentations make the model less sensitive to noise and less reliant on any particular textual cue. Finally, combining Prev/Prev augmentation further boosts translation performance (43.5 vs 42.7), as it reduces the reliance on previous GT sentences and better matches the inference setting.
LLM fine-tuning. In Tab. 2, we also examine a model variant without fine-tuning the LLM, but only training the mapping network with a frozen LLM. Comparing the last two rows, we observe that LoRA fine-tuning the decoder yields improvements across all metrics (B-RT: +2.9, IoU: +2.1, and LLM score: +0.29). We hypothesise that the improvement may partially be due to distinct linguistic characteristics of signed and spoken languages, but also due to adapting the LLM to our specific input structure.
| Model | B4 | B-RT | R-L | CIDEr | IoU | LLM |
| Albanie [2] | 1.0 | - | 10.2 | - | - | - |
| Sincan [65] | 1.3 | - | 8.9 | - | - | - |
| GFSLT [84] | 0.6 | 27.7 | 7.4 | 4.3 | 5.2 | 0.05 |
| Sign2GPT [75] | 0.7 | 34.3 | 10.6 | 12.8 | 8.2 | 0.37 |
| Sign2GPT (w/PGP) [75] | 0.9 | 35.2 | 11.4 | 16.1 | 8.7 | 0.49 |
| Ours (Vid+Prev+PG+BG) | 3.3 | 40.3 | 16.9 | 41.9 | 14.8 | 1.20 |
| Oracle | ||||||
| Sincan [65] (Vid+Prev) | 1.5 | 35.8 | 9.7 | 23.9 | 10.4 | 0.56 |
| Sincan [65] (Vid+Prev +Spot) | 2.9 | 37.0 | 12.4 | 41.0 | 12.5 | 0.80 |
| Ours (Vid+Prev+PG+BG) | 4.1 | 41.6 | 19.2 | 50.0 | 15.6 | 1.38 |
| Ours (Vid+Prev+Spot+BG) | 7.3 | 47.1 | 25.1 | 88.9 | 26.5 | 1.85 |
4.4 Comparison to the state of the art
BOBSL. We evaluate our model on the BOBSL test set, Sent-Test, using our full suite of evaluation metrics. We show in Tab. 3 that our approach gives a significant boost in all metrics compared to previous works. Specifically, compared to state-of-the-art approaches GFSLT [84] and Sign2GPT [75], we achieve a considerable increase in performance (40.3 vs 35.2 B-RT)—highlighting both the effectiveness of our method which leverages context and the increased challenge posed by the BOBSL dataset (as opposed to PHOENIX14T where [84, 75] were originally evaluated).
In the Oracle setup (the bottom block of Tab. 3), we compare to the setting of [65], where models have access to ground truth previous sentence and spottings extracted from the ground truth current sentence. When using the ground truth previous sentence at inference, our model outperforms [65] by a large margin (41.6 vs 35.8 B-RT). When replacing our pseudo-glosses with spottings, we further increase the margin, substantially outperforming their method (47.1 vs 37.0 B-RT, i.e. +10 point improvement).
How2Sign. Here, we demonstrate the generality of our method by training on the How2Sign dataset (see Sec. A.5 for details). In Tab. 4, we compare against the state of the art on the test set, and also report variants of our model by gradually adding more cues. We observe that adding the pseudo-glosses, as well as the contextual cue of the previous translated sentence boosts performance. We note that in this case, we do not use background descriptions since How2Sign does not consist of interpreted TV with an accompanying show. We find our best model (Vid+PG+Prev) achieves comparable performance with the state-of-the-art method VAP [36] in terms of B4 and attains a higher R-L score by nearly 5 points (32.5 vs 27.8). We note that we include the numbers from [82] and [61] (denoted with ), however, we do not compare to these as they train SLT additionally on a large ASL corpus of 984 hours [70].
| Model | B4 | B-RT | R-L | CIDEr | IoU | LLM |
| SSLT [82] | - | 55.7 | - | - | - | - |
| SSVP-SLT [61] | 15.5 | 49.6 | 38.4 | - | - | - |
| SSLT [82] | - | 34.0 | - | - | - | - |
| SSVP-SLT [61] | 7.0 | 39.3 | 25.7 | - | - | - |
| Fla-LLM [17] | 9.7 | - | 27.8 | - | - | - |
| VAP [36] | 12.9 | - | 27.8 | - | - | - |
| Ours (Vid) | 11.8 | 44.1 | 31.1 | 93.3 | 26.1 | 1.39 |
| Ours (Vid+PG) | 12.3 | 44.7 | 31.9 | 97.8 | 27.4 | 1.55 |
| Ours (Vid+PG+Prev) | 12.7 | 45.3 | 32.5 | 100.8 | 27.9 | 1.59 |
4.5 Qualitative analysis and limitations
We visually analyse how different input cues impact the translation outputs by providing relevant information beyond the signing video. In Fig. 3 (top left), a key focus of the sentence – the word jellyfish – is not signed. In Fig. 3 (top right), the sign for pronouns he and she is ambiguous. In such cases, the model needs to utilise available context – much like a human interpreter would – to accurately translate the sentence. By effectively leveraging the background context in both cases, the model is able to produce stronger translations. In other cases, context can be used to further augment information obtained from the video. This can be seen in Fig. 3 (middle left), where using only the video, the model gains a general theme about national forest but by leveraging the context, can precisely generate New Forest National Park. Fig. 3 (middle right) shows an example where the signer points to the background where the rhino horn appears on screen, and the similar sign for elephant that appears in pseudo-glosses is effectively suppressed.
However, we do observe several challenges as well: (i) different cues may present conflicting information, and while the model often learns to implicitly resolve such conflicts, there are cases where it struggles (see Fig. 3, bottom right); (ii) our model can struggle to discern the grammatical context of a sentence, e.g. it sometimes cannot distinguish whether a given sentence is a question or a statement; (iii) similarly the model makes frequent mistakes by missing negations; (iv) our model still faces difficulties with certain sign types, such as pointing and fingerspelling, which are essential components of sign language. These limitations highlight the complexity of sign language translation and underscore the need for continued research and development in the field. Additional qualitative results are provided in Appendix C.
5 Conclusion
In this work, we show that leveraging contextual information significantly enhances SLT performance in an open-vocabulary setting. Specifically, our framework utilises background descriptions from a captioning model and predictions of previous sentences, combined with pseudo-glosses and visual features. Through an extensive ablation study, we analyse the individual impact of each cue on sign language translation, and benchmark our method against previous state-of-the-art approaches on the BOBSL dataset to demonstrate its effectiveness. While our approach considerably outperforms prior methods, there remains a room for improvement in the translation quality to be deployed in real-world applications, necessitating to address the limitations mentioned above.
Acknowledgments. The images in this paper are used with kind permission of the BBC. This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011013395 made by GENCI. The authors would like to acknowledge the ANR project CorVis ANR-21-CE23-0003-01, a Google Research Scholar Award, and a Royal Society Research Professorship RPR1191132. The authors also thank Charles Raude, Prajwal KR, Prasanna Sridhar, Joon Son Chung, Makarand Tapaswi, Syrine Kalleli, and David Picard for their help.
APPENDIX
This appendix supplements the main paper by providing additional implementation details (Appendix A), experiments (Appendix B), and qualitative results (Appendix C).
[…]
Appendix A Implementation Details
We provide details on the LLM evaluation metric (Sec. A.1), the list of prompts for the translation model (Sec. A.2), the background description collation (Sec. A.3), the architectural design for the mapping network on BOBSL (Sec. A.4), and the training procedure on How2Sign (Sec. A.5).
A.1 LLM Evaluation metric
As mentioned in Sec. 4.1, our LLM-based evaluation metric is adapted from the CLAIR framework [14]. Here, we detail the prompts and show an analysis for this metric.
LLM evaluation prompt. Fig. A.1 shows the system, user, and assistant prompts, that we input to GPT-4o-mini [54], to define the sign language translation evaluation task. To calibrate the language model, we include 12 manually annotated in-context examples, displayed in Tab. A.9, with two examples per score from 0 to 5. Each example contains both the score and the reasoning according to our instructions, focusing on key nouns and verbs, while giving less importance to pronouns. This approach makes the metric interpretable, as the LLM outputs detailed reasoning for each score.
LLM evaluation analysis. As discussed in Sec. 4.3, we provide additional analysis and statistics on LLM-based evaluation. A human study was conducted where 5 annotators manually scored a set of 70 translations. Fig. A.2 and Fig. A.3 illustrate the correlation between the average of these human scores, various automatic metrics [56, 63, 45], and our LLM-based evaluation metric. As shown in Fig. A.2, the LLM-based metric exhibits the highest correlation with human judgments. This strong correlation highlights its potential as a useful method for evaluating sign language translations. We further provide qualitative examples for the LLM scores in Fig. A.4.
| Type | Prompt |
| Initial | You are an AI assistant designed to interpret a video of a sign language signing sequence and translate it into English. |
| Previous sentence | The previous context is the following: |
| Pseudo-glosses | The following are some possible words present in the sentence: |
| Background description | Description of the background is: |
| Visual features | The following are the video tokens: |
A.2 Prompt details
A.3 Background description collation
In Fig. A.5, we illustrate two examples to show the process for the background description collation. As explained in Sec. 3.2, in the first step, we extract captions from multiple frames; in the second step, we take the unique words after filtering out stop words.
We further perform several analyses on these background descriptions on the BOBSL training set. First, we measure the similarity between background descriptions and the ground truth translation sentences, and obtain 3.4% IoU, 5.3% precision, and 9.3% recall. We note that the informative signal in the background descriptions may be beyond the exact word overlap. Next, we look at the distribution of parts of speech, revealing 56.1% nouns, 19.4% verbs, 11.8% adjectives, and 7.9% proper nouns. Among the most frequently occurring words, “man” was identified as the most common noun, “standing” as the most common verb, and “front” as the most common adjective.
A.4 BOBSL mapping network
The details of the 2-layer MLP used as the mapping network are provided in Tab. A.2. The input to the mapping network consists of Video-Swin features, and its output serves as the input to the LLM decoder. Specifically, in our experiments, the size of the Video-Swin features is 768, while the input size of the LLM decoder (Llama3-8B) is 4,096.
| layer | input sizes | output sizes |
| fc1 | ||
| gelu | ||
| fc2 |
| layer | kernel | stride | padding | input sizes | output sizes |
| conv1 | 5 | 1 | 2 | ||
| relu1 | - | - | - | ||
| maxpool1 | 2 | 2 | - | ||
| conv2 | 5 | 1 | 2 | ||
| relu2 | - | - | - | ||
| maxpool2 | 2 | 2 | - | ||
| fc1 | - | - | - | ||
| gelu | - | - | - | ||
| fc2 | - | - | - |
A.5 How2Sign training details
ISLR training details. As mentioned in Sec. 3.2 of the main paper, we fine-tune the Video-Swin model, which is released by [60], with annotations provided by [23]. The training data is automatically annotated from mouthing and dictionary sources, and we set thresholds at 0.75 and 0.5, respectively, to filter the data and enhance its reliability. We train the model on 4 A6000 GPUs with a batch size of 24 per GPU, utilising the Adam optimizer [38]. Training is performed in bfloat16 precision. The training spans 30 epochs, including the warmup phase for the first 1 epoch. The learning rate is set to 0.0001 and one cycle cosine learning rate scheduler is adapted.
Visual features. We set the stride () to 1 for feature extraction using the Video-Swin model on the How2Sign dataset, as the data is smaller and manageable for training. The average number of features is 171, corresponding to a 6.8-second long sequence.
Mapping network on How2Sign. As mentioned in Sec. 3.3 of the main paper, we further provide detailed information about the mapping network for training our model on the How2Sign dataset. Through our experiments, we found that training with only 2-layer MLP was not successful on the How2Sign dataset. Therefore, we add a simple 1D CNN before the MLP layers to compress the long sequences with minimal additional parameters. The 1D CNN is configured with a specific sequence of layers: {K5, P2, K5, P2}, where denotes a kernel size of , and represents a pooling layer with a kernel size of [33]. Details of this mapping network, including input and output sizes, are provided in Tab. A.3.
| Model | Size | B4 | B-RT | R-L | CIDEr | IoU | LLM |
| Llama3.2 | 1B | 2.4 | 39.2 | 15.8 | 35.8 | 13.6 | 1.05 |
| Llama3.2 | 3B | 3.1 | 40.0 | 17.0 | 41.9 | 14.6 | 1.20 |
| Llama3 | 8B | 3.3 | 40.3 | 16.9 | 41.9 | 14.8 | 1.20 |
Target sentence augmentation for How2Sign. We observe overfitting starting from around 5–6 epochs when training with an LLM on the relatively small How2Sign dataset. To further improve the model’s performance, we employ a data augmentation technique that randomly drops 0–20% of the words from the GT sentences.
Appendix B Additional Experiments
We examine performance variations when using different LLM decoders (Sec. B.1), evaluate all possible cue combinations (Sec. B.2), and investigate scenarios with missing cues (Sec. B.3). We also showcase the performance of our ISLR backbone on the HowSign dataset (Sec. B.4), and report the reproduction results of GFSLT [84] and Sign2GPT [75] on the PHOENIX14T dataset (Sec. B.5).
B.1 Llama decoder variants
To further analyse the impact of the LLM decoder on performance, we experiment with various Llama variants. Specifically, we compare the Llama3-8B model used in the main paper experiments to more recent and smaller Llama3.2 models: Llama3.2-1B and Llama3.2-3B. As shown in Tab. A.4, Llama3.2-3B demonstrates performance comparable to Llama3-8B. When using the Llama3.2-1B model, we observe a performance drop of 0.7 in the BLEU-4 (B4) score compared to the Llama3-8B model. However, Llama3.2-1B still outperforms all baselines compared in Tab. 3 of the main paper. Note that this experiment is conducted on BOBSL Sent-Test.
| Vid | PG | Prev | BG | B-RT | IoU | LLM |
| ✓ | 41.0 | 16.6 | 1.29 | |||
| ✓ | ✓ | 41.8 | 17.5 | 1.40 | ||
| ✓ | ✓ | 41.5 | 17.0 | 1.38 | ||
| ✓ | ✓ | 41.9 | 17.5 | 1.41 | ||
| ✓ | ✓ | ✓ | 42.5 | 18.1 | 1.45 | |
| ✓ | ✓ | ✓ | 43.2 | 18.6 | 1.54 | |
| ✓ | ✓ | ✓ | 43.1 | 18.3 | 1.52 | |
| ✓ | ✓ | ✓ | ✓ | 43.5 | 18.8 | 1.56 |
B.2 Combining different cues
B.3 Missing cue scenario
As discussed in Sec. 3.3 of the main paper, the Drop Cue augmentation enables our model to perform flexible translations even when certain cues are missing during test time. The experimental results are presented in Tab. A.6. Note that, while the previous Tab. A.5 displays the performance of models trained with various cue combinations, Tab. A.6 reports the inference results of the model trained with all cues. Notably, our final model achieves results comparable to those of the models trained on specific combinations of cues (i.e. models listed in Tab. 1 of the main paper). This demonstrates that our final model can perform sign language translation with minimal performance degradation when certain cues are unavailable during inference.
| Vid | PG | Prev | BG | B-RT | IoU | LLM |
| ✓ | 41.1 | 17.2 | 1.31 | |||
| ✓ | ✓ | 41.8 | 18.1 | 1.41 | ||
| ✓ | ✓ | 41.7 | 17.3 | 1.39 | ||
| ✓ | ✓ | 42.0 | 17.0 | 1.42 | ||
| ✓ | ✓ | ✓ | 42.6 | 17.8 | 1.49 | |
| ✓ | ✓ | ✓ | 42.8 | 18.7 | 1.52 | |
| ✓ | ✓ | ✓ | 42.5 | 17.7 | 1.52 | |
| ✓ | ✓ | ✓ | ✓ | 43.5 | 18.8 | 1.56 |
| Model | Dataset | Per-instance | Per-class | ||
| top-1 | top-5 | top-1 | top-5 | ||
| I3D [23] | BOBSL How2Sign | 59.5 | 78.9 | 44.5 | 68.7 |
| Video-Swin (Ours) | How2Sign | 63.9 | 86.0 | 41.8 | 69.3 |
| Video-Swin (Ours) | BOBSL How2Sign | 77.0 | 92.8 | 58.5 | 82.3 |
B.4 ISLR performance on How2Sign
The test set provided by [23] is composed of 2,212 manually annotated data. We evaluate both per-instance and per-class accuracy metrics. Per-instance accuracy is calculated across all test instances, while per-class accuracy represents the average performance across the sign categories in the test set. This metric is particularly useful for addressing the unbalanced nature of the datasets, as recommended in [1].
As shown in Tab. A.7, our Video-Swin ISLR model, trained without pre-training on the BOBSL dataset, achieves performance comparable to the I3D ISLR model [23], which is pre-trained on the BOBSL dataset and fine-tuned on the How2Sign dataset. Furthermore, when the Video-Swin ISLR model is initialised with weights pre-trained on the BOBSL dataset, as released by [60], and further fine-tuned on the How2Sign dataset by us, it achieves a 13.1% improvement in per-instance top-1 accuracy and a 16.7% improvement in per-class top-1 accuracy. This underscores the effectiveness and robustness of our framework’s ISLR backbone.
| Model | B1 | B2 | B3 | B4 | Rouge |
| GFSLT [84] | 43.71 | 33.18 | 26.11 | 21.44 | 42.49 |
| GFSLT [84] | 42.02 | 31.88 | 25.30 | 20.76 | 42.62 |
| Sign2GPT [75] | 45.43 | 32.03 | 24.23 | 19.42 | 45.23 |
| Sign2GPT [75] | 44.14 | 32.72 | 25.49 | 20.82 | 43.70 |
| Sign2GPT (w/PGP) [75] | 49.54 | 35.96 | 28.83 | 22.52 | 48.90 |
| Sign2GPT (w/PGP) [75] | 46.90 | 35.72 | 28.30 | 23.22 | 46.28 |
B.5 Reproducing GFSLT and Sign2GPT on PHOENIX14T
As mentioned in Sec. 4.2 of the main paper, we reproduce the performance of the GFSLT and Sign2GPT models on the PHOENIX14T dataset [39, 11]. The results are shown in Tab. A.8. The symbol denotes the reproduced results, which show comparable performance to the results reported in their original papers across all metrics.
Training on BOBSL. For GFSLT, we observed that using the official codebase leads to gradient divergence during the masked word reconstruction process in text decoding. To mitigate this issue, we reduced the weight of the word reconstruction loss from 1 to 0.1. For Sign2GPT, as the official codebase only includes the model and hyperparameters, we developed training code using Accelerate [28] framework.
| 1 | Reference: | It’s blind to the genius loci. |
| Candidate: | And that’s what it means to be dislocated. | |
| Score: | 0 | |
| Reason: | No shared key nouns or verbs; the reference mentions ‘blind’ and ‘genius loci’, while the candidate mentions ‘dislocated’; meanings are different. | |
| 2 | Reference: | She put it by the entrance to the earth so we figure that they like heavy metal or something. |
| Candidate: | You’ve been in a wheelchair for a long time. | |
| Score: | 0 | |
| Reason: | No shared key nouns or verbs; the reference talks about ‘entrance’, ‘earth’, ‘heavy metal’, while the candidate mentions ‘wheelchair’; meanings are unrelated. | |
| 3 | Reference: | You’re coming along to the finale tomorrow? |
| Candidate: | I’ll have to wait until tomorrow. | |
| Score: | 1 | |
| Reason: | Shares the key noun ‘tomorrow’ but lacks other key content; meanings are somewhat related but differ. | |
| 4 | Reference: | A man’s can was open for attack at any point in their life. |
| Candidate: | It’s not a joke, it’s a way of life. | |
| Score: | 1 | |
| Reason: | Shares the key noun ‘life’, but overall meanings are different; reference discusses vulnerability, candidate discusses lifestyle. | |
| 5 | Reference: | Richard called English Nature, who told him that they were natterjack toads, Britain’s rarest amphibian. |
| Candidate: | Richard, a Nottinghamshire Englishman, is a naturalist. | |
| Score: | 2 | |
| Reason: | Shares key nouns ‘Richard’ and ‘English’; candidate omits details about ‘natterjack toads’ and ‘Britain’s rarest amphibian. | |
| 6 | Reference: | Cromwell treated Ireland like the primitive colony he thought it was, moving the Irish off their farms and using the land to pay his soldiers. |
| Candidate: | Cromwell was just one of many areas where the IRA set up roadblocks to stop loyalist paramilitaries and farmers from getting through. | |
| Score: | 2 | |
| Reason: | Shares key nouns ‘Cromwell’, ‘Ireland’, ‘farmers’; reference discusses historical actions, candidate discusses modern events; meanings differ. | |
| 7 | Reference: | He sort of guessed it would be 21 maybe 28 days, ended-up being 35. |
| Candidate: | He thought it was 21 days. | |
| Score: | 3 | |
| Reason: | Matches key nouns ‘he’, ‘thought’, ‘21 days’; candidate conveys a similar time estimation with fewer details. | |
| 8 | Reference: | They get as high as they can off the ground to push the other male down to the floor, and once that male becomes subservient, he slinks off and the dominant |
| male then goes to his female to breed. | ||
| Candidate: | But when one of them is selected, it’s the females that can be changed to breed with the new male. | |
| Score: | 3 | |
| Reason: | Shares key nouns ‘male’, ‘female’, ‘breed’; both discuss breeding behaviors, though specifics differ. | |
| 9 | Reference: | It’s a lack of understanding on both sides. |
| Candidate: | I don’t understand that. | |
| Score: | 3 | |
| Reason: | Shares key verb ‘understand’; both express lack of understanding; candidate is less specific. | |
| 10 | Reference: | OK, we’ll just ring him next time. |
| Candidate: | I’ll call you back. | |
| Score: | 4 | |
| Reason: | Shares key verb ‘call’ (synonym of ‘ring’); both involve making a call; minor differences in context. | |
| 11 | Reference: | Really excited. |
| Candidate: | I’m so excited. | |
| Score: | 5 | |
| Reason: | Conveys the same overall meaning; both express excitement with minor wording differences. | |
| 12 | Reference: | Every day is totally different. |
| Candidate: | You know, every day is different. | |
| Score: | 5 | |
| Reason: | Conveys the same overall meaning; both state that each day is different with minor phrasing differences. |
Appendix C Additional Qualitative Results
We present additional qualitative results similar to Fig. 3 of the main paper, where we display various inputs, and predictions from our final model compared to baselines. In the first sample of Fig. A.6, the previous sentence indirectly provides information related to location and area, allowing the model to successfully translate the word ‘river’. The second sample in Fig. A.6 demonstrates how the background description conveys information about the presence of multiple people on the screen. The third sample in Fig. A.6 demonstrates the ability of the background description to recognise characters on the screen. The first sample in Fig. A.7 shows the model accurately capturing the object of the sentence from the background description. The second sample in Fig. A.7 demonstrates that the models (Vid+PG, Vid+PG+Prev) make incorrect translations by referring to the ‘cliff’ word of the pseudo-glosses. However, when all cues are provided, the bias introduced by the pseudo-glosses is resolved. The third sample in Fig. A.7 shows a failure case where the model is biased by background information during the process of pronoun assignment.
Supplementary video. We provide a supplementary video on our project page showcasing several examples comparing against the baselines, along with dynamic sign language videos. The video further includes an example of continuous translation, where consecutive sign language sentences are translated.